OpenStudy (anonymous):

What is the sum of a 23–term arithmetic sequence where the first term is 9 and the last term is 185?

OpenStudy (anonymous):

sn=[n(a1+an) / 2] a1=9 a23=185 n=23 Plug into the sn equation

OpenStudy (anonymous):

Thanks ! Exactly what i needed.

OpenStudy (anonymous):

@DeeJay23 do you know what formula a_n has?

OpenStudy (anonymous):

uhmmm , Arthemetic Series ,i think ? im confused noww :/

OpenStudy (anonymous):

\[\LARGE a_n=a_1+(n-1)d\] but you don't have difference yet.. do you know how to find it?

OpenStudy (anonymous):

just substitute \[\LARGE a_{23}=a_1+(23-1)d\] \[\LARGE 185=9+22d\] solve for "d" then you have it. Once you're done with it... plug it back in the equation optiquest wrote: \[\LARGE S_n=\frac n2[2a_1+(n-1)d]\]